Use the sample data and confidence level given below to complete parts a through d.

A research institute poll asked respondents if they felt vulnerable to identity theft. In the poll, n=1040 and x=544 who said yes.

Use a 95% confidecne level.

A. find the best point of estimate of the population of portion p. B. Identify the value of the margin of error E. E= round to four decimal places as needed. C. Construct the confidence interval. _ < p <_ round to three decimal places. D. Write a statement that correctly interprets the confidence interval.
Answer #1

Given : n=1040 , X=544

a) The point estimate is ,

Point estimate=\hat{p}=\frac{X}{n}=\frac{544}{1040}=0.5231

b) Now , the critical value is ,

Za a/2 20.05/2 1.96 ; From Z-table

The margin of error is ,

E=Z_{\alpha/2}\sqrt{\frac{\hat{p}(1-\hat{p})}{n}}=1.96*\sqrt{\frac{0.5231(1-0.5231)}{1040}}=0.0304

c) The 95% confidence interval is ,

\hat{p}\pm E

0.5231\pm 0.0304

0.4927<p<0.5534